# Energy Conditions and Quantum Information

### v1

@@ 2.3 Geodesic congruences on curved spacetime
$\theta:=h^{\mu\nu}\nabla_{\nu}u_{\mu}=\nabla_{\mu}u^{\mu}\,,\quad\sigma_{\mu\nu}:=\nabla_{(\nu}u_{\mu)}-\dfrac{1}{3}\theta h_{\mu\nu}\,,\quad\omega_{\mu\nu}:=\nabla_{[\nu}u_{\mu]}\,$, (14)
@@
Let us consider a congruence of timelike geodesic $\gamma$ orthogonal to $3$-dimensional hypersurface $\Sigma$ with vanishing rotation $\omega_{\mu\nu}=0$. If we denote by $A$ an infinitesimal cross-sectional ($3$-dimensional) area element of the congruence, then
@@
$u^{\alpha}\nabla_{\alpha}\sigma_{\mu\nu}$ $=$ $-\frac{2}{3}\theta\sigma_{\mu\nu}-\left(\sigma_{\mu\sigma}\sigma_{\nu}{}^{\sigma}-\frac{1}{3}\sigma_{\alpha\beta}\sigma^{\alpha\beta}h_{\mu\nu}\right)$ (17)
$-C_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}+\frac{1}{2}\left(h_{\mu}{}^{\alpha}h_{\nu}{}^{\beta}-\frac{1}{3}h^{\alpha\beta}h_{\mu\nu}\right)R_{\alpha\beta}\,$,
@@
$C_{\mu\alpha\nu\beta}:=R_{\mu\alpha\nu\beta}+g_{\mu[\beta}R_{\nu]\alpha}+g_{\alpha[\nu}R_{\beta]\mu}+(1/3)Rg_{\mu[\nu}g_{\beta]\alpha}$ (18)
@@ 2.4 Focusing theorems and energy conditions
Now that we have formulas of the expansion and shear for both timelike and null geodesic congruence, we discuss the focusing effects. Below we consider first the null geodesic case in detail, and then derive essentially the same conclusion for the timelike case. We immediately see that the first and second terms in the right-hand side of (23) are non-positive definite, hence always tend to decrease the expansion $\theta$ and cause the convergence of the null geodesic congruence. The third term, $R_{\mu\nu}k^{\mu}k^{\nu}$, is non-positive if the NEC holds, thus also tends to decrease $\theta$ under the NEC. Consider the case that both the expansion and shear tensors vanish initially. If $R_{\mu\nu}k^{\mu}k^{\nu}\neq 0$ at some point of $\gamma$, then via the Raychaudhuri equation, $\theta$ becomes non-vanishing, which, again via the Raychaudhuri equation, makes $\theta$ start to decrease. Suppose now $R_{\mu\nu}k^{\mu}k^{\nu}=0$ (vacuum spacetime), and initially $\theta=0=\sigma_{\mu\nu}$. Even in this case, if the third term of equation (24) is non-vanishing, $p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0$, it makes the shear tensor non-vanishing by (24) and then produces, via (23), non-vanishing expansion. Then, again from (23), $\theta$ starts to decrease. Now we state the following condition. {itembox}[l]Definition 2.4.1 (Null generic condition) A spacetime is said to satisfy the null generic condition if every null geodesic (whose tangent vector field is denoted by $k^{\mu}$) possesses at least one point where
$k_{[\alpha}R_{\mu]\beta\nu[\gamma}k_{\sigma]}k^{\beta}k^{\nu}\neq 0$.
This condition is equivalent to
$\mbox{either (i) }\,R_{\mu\nu}k^{\mu}k^{\nu}\neq 0\,,\,\,\,\mbox{or (ii) }\,p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0\,$, (25)
@@
Suppose that our null geodesic curve $\gamma$ with the tangent vector $k^{\mu}$ is complete with respect to the affain parameter $\lambda$; namely $\gamma$ can be extended to take arbitrary value of $\lambda$ on the spacetime considered. Then, from the above observation, it is inevitable that under the generic conditions and the NEC, the expansion eventually negatively diverges $\theta\rightarrow$-$\infty$. In fact, under the NEC, we have from the Raychaudhuri equation (23),
@@
{itembox}
[l]Theorem 2.4.2 (Null focusing) Under the NEC, any hypersurface-orthogonal null geodesic congruence is focusing:
$\dfrac{{\rm d}\theta}{{\rm d}\lambda}\leqslant 0$.
Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\lambda=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the affine length
$\lambda\leqslant\dfrac{2}{|\theta_{0}|}$.
@@
Suppose the above null geodesic congruence containing $\gamma$ is emanating from a point $p\in\gamma$ at some $\lambda<0$, attains the negative value $\theta_{0}<0$ at $\lambda=0$, and as claimed above, reaches a point $q\in\gamma$ where $\theta\rightarrow$-$\infty$. Then, in general, $q$ is said to be conjugate to $p$. Then, one can claim that under the null generic condition and NEC, $\gamma$ must possess a pair of conjugate points (see, proposition 9.3.7 Wald). We can also consider a congruence of null geodesics with tangent vector field $k^{\mu}$ which are orthogonal to a $2$-dimensional spacelike surface $S$. There exist, in general, two families of such surface orthogonal null geodesics (often referred to as “ingoing” and “outgoing” null geodesics). In this case, a point $q$ away from $S$ along $\gamma$ where $\theta\rightarrow$-$\infty$ is said to be conjugate to $S$. The occurrence of a conjugate point $q\in\gamma$ to $p\in\gamma$ (or to $S$) is a natural consequence under the attractive nature of gravitation. An important and useful aspect of the occurrence of a conjugate point $q$ is that a null geodesic $\gamma$ fails to remain on the boundary of causal future (or past) of $p$ (or $S$). We shall discuss this aspect in more detail in the next subsection.
@@
{itembox}
[l]Definition 2.4.3 (Timelike generic condition) A spacetime is said to satisfy the timeilke generic condition if every timelike geodesic (whose tangent vector field is denoted by $u^{\mu}$) possesses at least one point where
$R_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}\neq 0$.
@@
{itembox}
[l]Theorem 2.4.4 (Timelike focusing) Under the SEC, any hypersurface-orthogonal timelike geodesic congruence is focusing:
$\dfrac{{\rm d}\theta}{{\rm d}\tau}\leqslant 0$.
Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\tau=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the proper time
$\tau\leqslant\dfrac{3}{|\theta_{0}|}$.
Similar to the null geodesic case, we can say that a point $q\in\gamma$ is conjugate to a point $p$ (or to $3$-dimensional hypersurface $\Sigma$) when the congruence is emanating from the point $p\in\gamma$ (or orthogonal to $\Sigma$). We can also claim that under the timeilke generic condition and the SEC, $\gamma$ must possess a pair of conjugate points. (See, proposition 9.3.1 Wald). An important aspect of the occurrence of a conjugate point $q$ for timelike geodesics is that a timelike geodesic $\gamma$ fails to be a local maximum of proper time between $q$ and $p$ (or $\Sigma$).
@@ 2.5 Singularity theorems and energy conditions
A spacetime singularity represents some extreme situation where general relativistic description of spacetime as a manifold breaks down due to infinite energy density or infinite spacetime curvature. This, in turn, implies that spacetime singularity should not be considered as part of the spacetime manifold. This leads to the idea that if a freely falling observer reaches a singularity, her/his world-line suddenly ends at a finite proper time and cannot be extended any further. Such a world-line (i.e., endless but never attains arbitrary large values in its proper time) can be described as an inextendible, incomplete timelike geodesic curve. The singularity theorems employ, as the criterion for the presence of a singularity, the existence of an inextendible incomplete causal (timelike or null) geodesic curve. For further motivation, satisfactory and unsatisfactory aspects of this definition of spacetime singularities, see, e.g., HE73 ; Wald84 .
@@
{itembox}
[l]Proposition 2.5.1 Any achronal null geodesic (hence, any achronal null hypersurface) does not admit a pair of conjugate points on it.
@@
Figure 3: Null geodesics orthogonally emanating from $S$. One such null geodesic $\gamma$ is passing through $p$, where its expansion changes sign from positive to negative, and then has a conjugate point $r$ in $J^{+}(p)$. Beyond this conjugate point, $\gamma$ fails to remain on the boundary $\partial J^{+}(S)$ and enters the interior $I^{+}(S)$. Since any point $q$ on $\gamma$ within $I^{+}(S)$ can be connected to $S$ by a timelike curve, $\gamma$ no longer is achronal beyond the conjugate point $r$. See also Figure 4 below.
An important example of achronal set is given by, {itembox}[l]Proposition 2.5.2 The boundary $\partial J^{+}(S)$ of the causal future of $S$ is a closed, $3$-dimensional, achronal submanifold.
Let $\Sigma$ be any closed achronal set. The future domain of dependence of $\Sigma$, denoted $D^{+}(\Sigma)$, is defined by the set of all points $p\in M$ such that every past-directed inextendible causal curve from $p$ intersects $\Sigma$. The past domain of dependence of $\Sigma$, denoted $D^{-}(\Sigma)$ is defined analogously. One often considers the union $D(\Sigma):=D^{+}(\Sigma)\cup D^{-}(\Sigma)$. In general, when $D(\Sigma)$ is a proper subset of $M$, i.e., $D(\Sigma)\varsubsetneqq M$, the boundary $\partial D^{\pm}(\Sigma)$ is called the future (past) Cauchy horizon of $\Sigma$, denoted by $H^{\pm}(\Sigma)$, respectively. A physical importance of the notion of the domain of dependence is that the physics inside $D(\Sigma)$ can be determined by the initial data on $\Sigma$, without any other information, e.g., boundary conditions outside $D(\Sigma)$.
@@
Then, one has {itembox}[l]Proposition 2.5.3 (Existence of maximum length curve) Let $N$ be globally hyperbolic (possibly as a subset of $M$). Then, for any two points $p,q\in N$ with $q\in J^{+}(p)$, there exists a causal geodesic curve $\mu$ from $p$ to $q$ which attains the maximum length among all causal curves from $p$ to $q$.
@@
As mentioned briefly just below the focusing theorems in the previous subsection, if a timelike geodesic connecting two points $p,q\in M$ contains a pair of conjugate points between the two, then there exists another causal curve whose length (proper time) is larger than the original timelike geodesic. The Lorentzian distance function $d:M\times M\rightarrow{\mathbb{R}}\cup\infty$ between two points $p,q\in M$ is defined by $d(p,q):=\sup\{L(\lambda):\lambda\in C(p,q)\}$. If $p,q\in I^{+}(p),\>p\neq q$, then $d(p,q)>0$, and if $q\in E^{+}(p)$ or not causally related, then $d(p,q)=0$. For any $q,r\in I^{+}(p),\>q\in I^{+}(r)$,
@@
Similarly, if a null geodesic connecting two points $p,q\in M$ contains a pair of conjugate points on it, then it cannot be achronal since $p,q$ can be then connected by some timelike curve. In other words, it fails to be the fastest causal curve. Combining this observation with Prop. 2.5.3, we have the following:
{itembox}
[l]Proposition 2.5.4 Any causal geodesic curve which connects two points $p$ and $q$, and which attains the maximum length between them cannot have a pair of conjugate points between $p$ and $q$.
@@
    The first singularity theorem by Penrose Penrose:1964wq assumes the global hyperbolicity, or equivalently the existence of a Cauchy surface, as a condition for global causal structure (Penrose’s theorem actually requires a non-compact Cauchy surface). However, the global hyperbolicity may be considered to be too strong a causality condition. After demonstrating contradictory consequences, instead of claiming that the assumption of the geodesic completeness fails, one may rather be able to conclude that the assumption of the global hyperbolicity must fail. In this respect, it is worth noting that under the same conditions of Penrose’s theorem, except the assumption of the global hyperbolicity, one can construct geodesically complete spacetimes, among which most well-known example is the Bardeen’s regular black hole.
    For this reason, after Penrose’s theorem Hawking:1970zqf , Hawking and Penrose proposed the improved version of the singularity theorem, in which the unwanted assumption of the global hyperbolicity is dropped, and instead, the chronology condition—which demands that no closed timelike curves exist—is imposed. This is a reasonable condition as to consider a physically realistic universe.
@@
    This property is described by the requirement of the trapped set defined as follows: {itembox}[l]Definition 2.5.5 (Future trapped set) A closed achronal subset $S$ is said to be a future trapped set if $E^{+}(S)$ is compact. The past trapped set is defined analogously. A simple example is a $2$-dimensional closed surface $S$ in the static Einstein universe $M={\mathbb{R}}\times S^{3}$ with the metric,
@@
    Figure 5: An achronal subset $S$ in spatially compact universe with $\Sigma$ being a compact Cauchy surface. The null geodesics (thick curves) generating the future horismos $E^{+}(S)$ have the two future endpoints due to the compactness of the universe, and accordingly $E^{+}(S)$ also is compact, thus $S$ is a trapped set. One may think of $S$ as any $2$-sphere of constant $T$ and $R$ in the static Einstein universe. (The two points labeled as $S$ in the figure represent the single $2$-sphere $S$.) $E^{+}(S)$ has the two future endpoints; one for the “ingoing” null geodesics, and the other for the “outgoing” ones, which are orthogonal to $S$. It is important to note, however, that in a closed universe, one cannot generally distinguish between ingoing and outgoing.
    In the context of gravitational collapse and black hole physics, the most important example of a trapped set is the closed trapped surface, which is a compact $2$-dimensional surface $\cal T$} such that the expansions $\theta_{\pm}$ of the two families of future (or past) null geodesic congruence orthogonal to $\cal T$} both are negative. An example is any $2$-sphere of constant area radius inside the Schwarzschild (horizon) radius $r_{h}:=2G{\cal M}$. To be concrete, let us express the Schwarzschild metric666 In the standard coordinates $x^{\mu}=(t,r,\vartheta,\varphi)$, the Schwarzschild metric is given as ${\rm d}s^{2}=-f(r)\,{\rm d}t^{2}+f^{-1}(r)\,{\rm d}r^{2}+r^{2}({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (33) where $f(r):=1-2G{\cal M}/r$, ${\cal M}=const.$. In one of the exterior region, U<0,V>0U<0,\>V>0, the two coordinates are related as $U=-e^{-(t-r_{*})/2r_{h}},\>V=e^{(t+r_{*})/2r_{h}}$, where $r_{*}=r+r_{h}\log(r-r_{h})$. in the Kruskal coordinates $\bar{x}^{\mu}=(U,V,\vartheta,\varphi)$,
@@
Let us state Penrose’s version of the singularity theorem Penrose:1964wq . {itembox}[l]Theorem 2.5.6 (Penrose 1965) Spacetime $(M,g_{\mu\nu})$ cannot be null geodesically complete if:
  • (i)
    there is non-compact Cauchy surface $\Sigma$ in $M$,
  • (ii)
    there is a closed trapped surface $\cal T$} in $M$,
  • (iii)
    the NEC holds.
@@
Next, consider Hawking and Penrose’s improved version of the theorem Hawking:1970zqf , in which the assumption of the global hyperbolicity (i.e., the existence of a Cauchy surface) is eliminated: {itembox}[l]Theorem 2.5.7 (Hawking-Penrose 1970) The following three conditions cannot all hold:
  • (i)
    the chronology condition holds,
  • (ii)
    there is a trapped set $\cal S$},
  • (iii)
    every inextendible causal geodesic contains a pair of conjugate points.

### v2

@@ 2.3 Geodesic congruences on curved spacetime
$\theta:=h^{\mu\nu}\nabla_{\nu}u_{\mu}=\nabla_{\mu}u^{\mu}\,,\quad\sigma_{\mu\nu}:=\nabla_{(\nu}u_{\mu)}-\dfrac{1}{3}\theta h_{\mu\nu}\,$,
$\omega_{\mu\nu}:=\nabla_{[\nu}u_{\mu]}\,$, (14)
@@
Let us consider a congruence of timelike geodesics orthogonal to $3$-dimensional hypersurface $\Sigma$ with vanishing rotation $\omega_{\mu\nu}=0$. If we denote by $A$ an infinitesimal cross-sectional ($3$-dimensional) area element of the congruence, then
@@
$u^{\alpha}\nabla_{\alpha}\sigma_{\mu\nu}$ $=$ $-\frac{2}{3}\theta\sigma_{\mu\nu}-\left(\sigma_{\mu\sigma}\sigma_{\nu}{}^{\sigma}-\frac{1}{3}\sigma_{\alpha\beta}\sigma^{\alpha\beta}h_{\mu\nu}\right)-C_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}+\frac{1}{2}\left(h_{\mu}{}^{\alpha}h_{\nu}{}^{\beta}-\frac{1}{3}h^{\alpha\beta}h_{\mu\nu}\right)R_{\alpha\beta}\,$, (17)
@@
$C_{\mu\alpha\nu\beta}$ $:=$ $R_{\mu\alpha\nu\beta}+g_{\mu[\beta}R_{\nu]\alpha}+g_{\alpha[\nu}R_{\beta]\mu}+\dfrac{1}{3}Rg_{\mu[\nu}g_{\beta]\alpha}$ (18)
@@ 2.4 Focusing theorems and energy conditions
Now that we have formulas of the expansion and shear for both timelike and null geodesic congruence, we discuss the focusing effects. Below we consider first the null geodesic case in detail, and then derive essentially the same conclusion for the timelike case. We immediately see that the first and second terms in the right-hand side of (23) are non-positive definite, hence always tend to decrease the expansion $\theta$ and cause the convergence of the null geodesic congruence. The third term, $R_{\mu\nu}k^{\mu}k^{\nu}$, is non-positive if the NEC holds, thus also tends to decrease $\theta$ under the NEC. Consider the case that both the expansion and shear tensors vanish initially. If $R_{\mu\nu}k^{\mu}k^{\nu}\neq 0$ at some point of $\gamma$, then via the Raychaudhuri equation, $\theta$ becomes non-vanishing, which, again via the Raychaudhuri equation, makes $\theta$ start to decrease. Suppose now $R_{\mu\nu}k^{\mu}k^{\nu}=0$ (vacuum spacetime), and initially $\theta=0=\sigma_{\mu\nu}$. Even in this case, if the third term of equation (24) is non-vanishing, $p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0$, it makes the shear tensor non-vanishing by (24) and then produces, via (23), non-vanishing expansion. Then, again from (23), $\theta$ starts to decrease. Now we state the following condition.
  Definition 2.4.1 (Null generic condition)   A spacetime is said to satisfy the null generic condition if every null geodesic (whose tangent vector field is denoted by $k^{\mu}$) possesses at least one point where $k_{[\alpha}R_{\mu]\beta\nu[\gamma}k_{\sigma]}k^{\beta}k^{\nu}\neq 0$.     This condition is equivalent to either (i) or (ii) below holds Wald84 ,
$\mbox{(i) }\,R_{\mu\nu}k^{\mu}k^{\nu}\neq 0\,,\,\,\mbox{(ii) }\,p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0\,$, (25)
@@
Suppose that our null geodesic curve $\gamma$ with the tangent vector $k^{\mu}$ is complete with respect to the affine parameter $\lambda$; namely $\gamma$ can be extended to take arbitrary value of $\lambda$ on the spacetime considered. Then, from the above observation, it is inevitable that under the generic conditions and the NEC, the expansion eventually negatively diverges $\theta\rightarrow$-$\infty$. In fact, under the NEC, we have from the Raychaudhuri equation (23),
@@
  Theorem 2.4.2 (Null focusing)   Under the NEC, any hypersurface-orthogonal null geodesic congruence is focusing: $\dfrac{{\rm d}\theta}{{\rm d}\lambda}\leqslant 0$. Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\lambda=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the affine length $\lambda\leqslant\dfrac{2}{|\theta_{0}|}$.
@@
Suppose the above null geodesic congruence containing $\gamma$ is emanating from a point $p\in\gamma$ at some $\lambda<0$, attains the negative value $\theta_{0}<0$ at $\lambda=0$, and as claimed above, reaches a point $q\in\gamma$ where $\theta\rightarrow$-$\infty$. Then, in general, $q$ is said to be conjugate to $p$. Then, one can claim that under the null generic condition and NEC, $\gamma$ must possess a pair of conjugate points (see, proposition 9.3.7 Wald Wald84 ). We can also consider a congruence of null geodesics with tangent vector field $k^{\mu}$ which are orthogonal to a $2$-dimensional spacelike surface $S$. There exist, in general, two families of such surface orthogonal null geodesics (often referred to as “ingoing” and “outgoing” null geodesics). In this case, a point $q$ away from $S$ along $\gamma$ where $\theta\rightarrow$-$\infty$ is said to be conjugate to $S$. The occurrence of a conjugate point $q\in\gamma$ to $p\in\gamma$ (or to $S$) is a natural consequence under the attractive nature of gravitation. An important and useful aspect of the occurrence of a conjugate point $q$ is that a null geodesic $\gamma$ fails to remain on the boundary of causal future (or past) of $p$ (or $S$). We shall discuss this aspect in more detail in the next subsection.
@@
  Definition 2.4.3 (Timelike generic condition)   A spacetime is said to satisfy the timelike generic condition if every timelike geodesic (whose tangent vector field is denoted by $u^{\mu}$) possesses at least one point where $R_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}\neq 0$.
@@
  Theorem 2.4.4 (Timelike focusing)   Under the SEC, any hypersurface-orthogonal timelike geodesic congruence is focusing: $\dfrac{{\rm d}\theta}{{\rm d}\tau}\leqslant 0$. Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\tau=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the proper time $\tau\leqslant\dfrac{3}{|\theta_{0}|}$.
Similar to the null geodesic case, we can say that a point $q\in\gamma$ is conjugate to a point $p$ (or to $3$-dimensional hypersurface $\Sigma$) when the congruence is emanating from the point $p\in\gamma$ (or orthogonal to $\Sigma$). We can also claim that under the timelike generic condition and the SEC, $\gamma$ must possess a pair of conjugate points. (See, proposition 9.3.1 Wald Wald84 ). An important aspect of the occurrence of a conjugate point $q$ for timelike geodesics is that a timelike geodesic $\gamma$ fails to be a local maximum of proper time between the initial point $p$ (or $\Sigma$) and any point beyond $q$ HE73 ; Wald84 .
@@ 2.5 Singularity theorems and energy conditions
A spacetime singularity represents some extreme situation where the general relativistic description of spacetime as a manifold breaks down due to infinite energy density or infinite spacetime curvature. This, in turn, implies that spacetime singularity should not be considered as part of the spacetime manifold. This leads to the idea that if a freely falling observer reaches a singularity, her/his world-line cannot be continued beyond a certain proper time. Such a world-line (i.e., endless but never attains arbitrary large values in its proper time) can be described as an inextendible, incomplete timelike geodesic curve. The singularity theorems employ, as the criterion for the presence of a singularity, the existence of an inextendible incomplete causal (timelike or null) geodesic curve. For further motivation, satisfactory and unsatisfactory aspects of this definition of spacetime singularities, see, e.g., HE73 ; Wald84 .
@@
  Proposition 2.5.1   Any achronal null geodesic (hence, any achronal null hypersurface) does not admit a pair of conjugate points on it.
@@
Figure 3: Null geodesics emanating from a subset $S$ of $M$. One such null geodesic $\gamma$ is passing through $p$, where its expansion changes sign from positive to negative, and then has a conjugate point $r$ in $J^{+}(p)$. Beyond this conjugate point, $\gamma$ fails to remain on the boundary $\partial J^{+}(S)$ and enters the interior $I^{+}(S)$. Since any point $q$ on $\gamma$ within $I^{+}(S)$ can be connected to $S$ by a timelike curve, $\gamma$ no longer is achronal beyond the conjugate point $r$. See also Figure 4 below.
An important example of achronal set is given by,
  Proposition 2.5.2   The boundary $\partial J^{+}(S)$ of the causal future of $S$ is a closed, $3$-dimensional, achronal submanifold.
Let $M$ be connected and let $\Sigma$ be any closed achronal set in $M$. The future domain of dependence of $\Sigma$, denoted $D^{+}(\Sigma)$, is defined by the set of all points $p\in M$ such that every past-directed inextendible causal curve from $p$ intersects $\Sigma$. The past domain of dependence of $\Sigma$, denoted $D^{-}(\Sigma)$ is defined analogously. One often considers the union $D(\Sigma):=D^{+}(\Sigma)\cup D^{-}(\Sigma)$. In general, when $D(\Sigma)$ is a proper subset of $M$, i.e., $D(\Sigma)\varsubsetneqq M$, the boundary $\partial D^{\pm}(\Sigma)$ is called the future (past) Cauchy horizon of $\Sigma$, denoted by $H^{\pm}(\Sigma)$, respectively. A physical importance of the notion of the domain of dependence is that the physics inside $D(\Sigma)$ can be determined by the initial data on $\Sigma$, without any other information, e.g., boundary conditions outside $D(\Sigma)$.
@@
Then, one has
  Proposition 2.5.3 (Maximum length curve)   Let $N$ be globally hyperbolic (possibly as a subset of $M$). Then, for any two points $p,q\in N$ with $q\in J^{+}(p)$, there exists a causal geodesic curve $\mu$ from $p$ to $q$ which attains the maximum length among all causal curves from $p$ to $q$.
@@
As mentioned briefly just below the focusing theorems in the previous subsection, if a timelike geodesic connecting two points $p,q\in M$ contains a pair of conjugate points between the two, then there exists another causal curve whose length (proper time) is larger than the original timelike geodesic. The Lorentzian distance function $d:M\times M\rightarrow{\mathbb{R}}\cup\{\infty\}$ between two points $p,q\in M$ is defined by $d(p,q):=\sup\{L(\lambda):\lambda\in C(p,q)\}$. If $p,q\in I^{+}(p),\>p\neq q$, then $d(p,q)>0$, and if $q\in E^{+}(p)$ or not causally related, then $d(p,q)=0$. For any $q,r\in I^{+}(p),\>q\in I^{+}(r)$,
@@
Similarly, if a null geodesic connecting two points $p,q\in M$ contains a pair of conjugate points on it, then it cannot be achronal since $p,q$ can be then connected by some timelike curve. Combining this observation with Prop. 2.5.3, we have the following:
  Proposition 2.5.4   Any causal geodesic curve which connects two points $p$ and $q$, and which attains the maximum length between them cannot have a pair of conjugate points between $p$ and $q$.
@@
    The first singularity theorem by Penrose Penrose:1964wq assumes the global hyperbolicity, or equivalently the existence of a Cauchy surface, as a condition for global causal structure (Penrose’s theorem actually requires a non-compact Cauchy surface). However, the global hyperbolicity may be considered to be too strong a causality condition. After demonstrating contradictory consequences, instead of claiming that the assumption of the geodesic completeness fails, one may rather be able to conclude that the assumption of the global hyperbolicity must fail. In this respect, it is worth noting that under the same conditions of Penrose’s theorem, except the assumption of the global hyperbolicity, one can construct geodesically complete spacetimes, among which most well-known example is the Bardeen’s regular black hole 2767662 .
    For this reason, after Penrose’s theorem Penrose:1964wq , Hawking and Penrose Hawking:1970zqf proposed the improved version of the singularity theorem, in which the unwanted assumption of the global hyperbolicity is dropped, and instead, the chronology condition—which demands that no closed timelike curves exist—is imposed. This is a reasonable condition as to consider a physically realistic universe.
@@
    This property is described by the requirement of the trapped set defined as follows:
      Definition 2.5.5 (Future trapped set)   A closed achronal subset $S$ is said to be a future trapped set if $E^{+}(S)$ is compact.     The past trapped set is defined analogously. A simple example is a $2$-dimensional closed surface $S$ in the static Einstein universe $M={\mathbb{R}}\times S^{3}$ with the metric,
@@
    Figure 5: An achronal subset $S$ in spatially compact universe with $\Sigma$ being a compact Cauchy surface. The null geodesics (thick curves) generating the future horismos $E^{+}(S)$ have the two future endpoints due to the compactness of the universe, and accordingly $E^{+}(S)$ also is compact, thus $S$ is a trapped set. One may think of $S$ as any $2$-sphere of constant $T$ and $R$ in the static Einstein universe. (The two points labeled as $S$ in the figure represent the single $2$-sphere $S$.) $E^{+}(S)$ has the two future endpoints; one for the “ingoing” null geodesics, and the other for the “outgoing” ones, which are orthogonal to $S$. It is important to note, however, that in a closed universe, one cannot generally distinguish between ingoing and outgoing. Note also that in the static Einstein universe, every Cauchy surface $\Sigma$ itself is in fact a trapped set since $E^{+}(\Sigma)=\Sigma$ is compact.
    In the context of gravitational collapse and black hole physics, the most important example of a trapped set is the closed trapped surface, which is a compact, orientable, and spacelike $2$-dimensional surface $\cal T$} such that the expansions $\theta_{\pm}$ of the two families of future (or past) null geodesic congruence orthogonal to $\cal T$} both are negative. An example is any $2$-sphere of constant area radius inside the Schwarzschild (horizon) radius $r_{h}:=2G{\cal M}$. To be concrete, let us express the Schwarzschild metric [^7] In the standard coordinates $x^{\mu}=(t,r,\vartheta,\varphi)$, the Schwarzschild metric is given as ${\rm d}s^{2}=-f(r)\,{\rm d}t^{2}+f^{-1}(r)\,{\rm d}r^{2}+r^{2}({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (33) where $f(r):=1-2G{\cal M}/r$, ${\cal M}=const.$. In one of the exterior region, U<0,V>0U<0,\>V>0, the two coordinates are related as $U=-e^{-(t-r_{*})/2r_{h}},\>V=e^{(t+r_{*})/2r_{h}}$, where $r_{*}=r+r_{h}\log(r-r_{h})$. in the Kruskal coordinates $\bar{x}^{\mu}=(U,V,\vartheta,\varphi)$,
@@
Let us state Penrose’s version of the singularity theorem Penrose:1964wq .
  Theorem 2.5.6 (Penrose 1965)   Spacetime $(M,g_{\mu\nu})$ cannot be null geodesically complete if: (i) there is non-compact Cauchy surface $\Sigma$ in $M$, (ii) there is a closed trapped surface $\cal T$} in $M$, (iii) the NEC holds.
@@
Next, consider Hawking and Penrose’s improved version of the theorem Hawking:1970zqf , in which the assumption of the global hyperbolicity (i.e., the existence of a Cauchy surface) is eliminated:
  Theorem 2.5.7 (Hawking-Penrose 1970)   The following three conditions cannot all hold: (i) the chronology condition holds, (ii) there is a trapped set $\cal S$}, (iii) every inextendible causal geodesic contains a pair of conjugate points.